OpenStudy (anonymous):

if the equation x^2-kx-36=0 has x=12 as one root what is the value of k?

OpenStudy (anonymous):

That would mean f(12) = 0. (12)² - 12k - 36 = 0 144 - 12k - 36 = 0 -12k = -108 k = 9

OpenStudy (mathstudent55):

Another way of solving this problem is: If x = 12 is a root, then x^2 - kx - 36 has to factor into (x - 12)(x + 3) since you need the 3 to multiply the -12 to give you -36. (x - 12)(x + 3) = x^2 + 3x - 12x - 36 = x^2 - 9x - 36 When you compare x^2 - 9x - 36 with x^2 - kx - 36, you get k = 9.